# MCDFN: Supply Chain Demand Forecasting via an Explainable Multi-Channel Data Fusion Network Model Integrating CNN, LSTM, and GRU

### v1

@@ Abstract
Accurate demand forecasting is critical for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, often based on simple arithmetic, fail to capture the complex patterns introduced by seasonal variability and special events. Despite the advancements in deep-learning techniques, the absence of interpretable forecasting models poses a significant challenge in understanding and explaining model predictions. In response, we introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture designed to integrate multiple data modalities for superior demand forecasting. MCDFN combines Convolutional Neural Networks (CNN), Long Short-Term Memory networks (LSTM), and Gated Recurrent Units (GRU) to extract spatial and temporal features from input time series data, thereby enhancing predictive performance. Our research rigorously validates MCDFN’s efficacy through extensive benchmarking against seven other deep-learning models. Notably, MCDFN outperforms its counterparts across key metrics, boasting impressive MSE (23.5738%), RMSE (4.8553%), MAE (3.9991%), and MAPE (20.1575%) scores. Additionally, MCDFN’s predictions were shown to have no significant difference from actual values through a paired t-test with a p-value of 5%. A 10-fold cross-validated statistical paired t-test further confirmed MCDFN’s robustness, significantly outperforming other models at a 5% level of significance. To address the inherent “black box” nature of MCDFN, we employ explainable artificial intelligence (XAI) techniques like ShapTime and Permutation Feature Importance, providing insights into model decision-making processes. Our research advances demand forecasting methodologies and provides practical implementation guidelines for seamlessly integrating MCDFN into existing supply chain systems. Moreover, we outline future research directions to enhance MCDFN’s scalability, integrate external factors for improved predictive capabilities, and develop user-friendly interfaces for real-world deployment.
@@
Supply Chain Management , Demand Forecasting , Deep Learning , CNN , LSTM , GRU
journal: Expert Systems with Applications
@@ 1 Introduction
  8. 8.
    We provide detailed implementation guidelines and a comprehensive framework for integrating MCDFN into existing SCM systems, ensuring ease of adoption and practical relevance for industry professionals.
@@ 3.2.8 Proposed MCDFN Model
We used hyperparameter tuning to optimize the number of filters in the CNN layers, kernel sizes, and units in the LSTM and GRU layers. After 90 trials of hyperparameter search, the optimal CNN filters, kernel size, and LSTM/GRU units were found to be 352, 1, and 64, respectively, in the 16th trial. The model with these hyperparameters achieved the lowest validation MSE of 23.29% during trials and tuning process.

### v2

@@ Abstract
Accurate demand forecasting is vital for optimizing supply chain management and enhancing organizational resilience. Traditional forecasting methods, relying on simple arithmetic, often fail to capture complex patterns caused by seasonal variability and special events. Although deep learning techniques have advanced, the lack of interpretable models hampers understanding and explaining predictions. We introduce the Multi-Channel Data Fusion Network (MCDFN), a novel hybrid deep learning architecture integrating multiple data modalities for superior demand forecasting. MCDFN utilizes Convolutional Neural Networks, Long Short-Term Memory networks, and Gated Recurrent Units to extract spatial and temporal features from time series data. Extensive benchmarking against seven other deep-learning models validates MCDFN’s efficacy, showing it outperforms its counterparts across key metrics with a mean squared error of 23.5738%, root mean squared error of 4.8553%, mean absolute error of 3.9991%, and mean absolute percentage error of 20.1575%. A paired t-test with a p-value of 5% indicated no significant difference between MCDFN’s predictions and actual values. A 10-fold cross-validated statistical paired t-test further confirmed its robustness, significantly outperforming other models at a 5% significance level. To address the ”black box” nature of MCDFN, we employ explainable AI techniques such as ShapTime and Permutation Feature Importance, offering insights into model decision-making processes. This research advances demand forecasting methodologies and provides practical guidelines for integrating MCDFN into existing supply chain systems while also outlining future research directions to enhance scalability, integrate external factors for better predictions, and develop user-friendly interfaces for real-world applications.
@@
Supply Chain Demand Forecasting , Explainable Artificial Intelligence , Deep Learning , Convolutional Neural Networks , Long Short-Term Memory , Gated Recurrent Units
@@ 3.2.8 Proposed MCDFN Model
We used hyperparameter tuning to optimize the number of filters in the CNN layers, kernel sizes, and units in the LSTM and GRU layers. After 90 trials of hyperparameter search, the optimal CNN filters, kernel size, and LSTM/GRU units were found to be 352, 1, and 64, respectively, in the 16th trial. The model with these hyperparameters achieved the lowest validation MSE of 23.29% during trials and the tuning process.